### v1

# MambaIO: Global-Coordinate Inertial Odometry for Pedestrians via Multi-Scale Frequency-Decoupled Modeling

Abstract
Inertial Odometry (IO) enables real-time localization using only acceleration and angular velocity measurements from an Inertial Measurement Unit (IMU), making it a promising solution for localization in consumer-grade applications. Traditionally, IMU measurements in IO have been processed under two coordinate system paradigms: the body coordinate frame and the global coordinate frame, with the latter being widely adopted. However, recent studies in drone scenarios have demonstrated that the body frame can significantly improve localization accuracy, prompting a re-evaluation of the suitability of the global frame for pedestrian IO. To address this issue, this paper systematically evaluates the effectiveness of the global coordinate frame in pedestrian IO through theoretical analysis, qualitative inspection, and quantitative experiments. Building upon these findings, we further propose MambaIO, which decomposes IMU measurements into high-frequency and low-frequency components using a Laplacian pyramid. The low-frequency component is processed by a Mamba architecture to extract implicit contextual motion cues, while the high-frequency component is handled by a convolutional structure to capture fine-grained local motion details. Experiments on multiple public datasets show that MambaIO substantially reduces localization error and achieves state-of-the-art (SOTA) performance. To the best of our knowledge, this is the first application of the Mamba architecture to the inertial odometry task.
I Introduction
The Inertial Measurement Unit (IMU) serves as the primary data source for Inertial Odometry (IO), providing acceleration and angular velocity measurements[surveyILS]. Owing to its low cost, compact size, and minimal power consumption, the IMU is widely integrated into mobile devices such as smartphones, robots, and unmanned aerial vehicles (UAV).
IO aims to estimate the trajectory of a moving carrier (e.g., humans or drones) based solely on acceleration and angular velocity measurements from the IMU, enabling high-precision self-localization and tracking. Traditional IO approaches primarily rely on numerical integration, but noise and bias inherent in IMU measurements accumulate over time, leading to significant drift errors. In contrast, learning-based IO methods, trained end-to-end on large-scale IMU datasets, automatically uncover latent motion regularities and demonstrate superior performance in terms of localization accuracy, robustness, and cross-scenario generalizability[RoNIN].
Early learning-based IO methods attempted to estimate the pose of the carrier directly in the body coordinate frame. However, RoNIN reported that the body frame undergoes frequent changes during motion, resulting in discontinuous motion representations and associated issues. To mitigate these problems, RoNIN introduced a Heading-Agnostic Coordinate Frame (HACF), defined as any coordinate system whose $Z$-axis is aligned with the gravity direction, thereby providing a more stable and reliable motion representation. This approach has since become the prevailing data processing paradigm in the IO community. Nevertheless, recent studies in drone-based scenarios have shown that, under certain conditions, using the body frame directly can lead to substantially improved localization accuracy. This counterintuitive observation motivates a reassessment of a fundamental question:
Is the global coordinate representation truly beneficial for pedestrian IO?
To address this question, we first conduct a systematic analysis of the representational characteristics of inertial data under different coordinate frames. Through kinematic analysis and visualization, we verify that the global coordinate frame provides more discriminative feature representations for pedestrian motion. Building on this finding, we propose MambaIO, which decomposes IMU measurements into frequency components to enable targeted learning. Experimental results demonstrate that MambaIO further improves localization accuracy in pedestrian IO, with selected results shown in Fig. 3.
The main contributions of this work are summarized as follows:
  • Coordinate Frame Representation Analysis: We systematically compare the representational capabilities of the global and body-fixed coordinate frames and their impact on learning performance, empirically validating the necessity of adopting the global frame in pedestrian IO.
  • Laplacian Pyramid Decomposition: We propose a PyTorch-based Laplacian pyramid method that efficiently decomposes IMU measurements into high- and low-frequency components, thereby enhancing the specificity and effectiveness of subsequent feature learning.
  • Mamba Module: We introduce the Mamba architecture to specifically model the contextual motion information embedded in the low-frequency IMU measurements component.
  • Multi-Path Convolutional Module: We design a multi-scale convolutional module to extract local motion features from the high-frequency IMU measurements component, which collaborates with the Mamba module to achieve comprehensive modeling of motion dynamics.
II Related Work
From a methodological perspective, research on IO can be broadly divided into two categories: traditional methods based on Newtonian mechanics and data-driven approaches leveraging machine learning.
Newtonian Mechanics-Based Methods: Early approaches, exemplified by Strapdown Inertial Navigation Systems (SINS), estimate the position and orientation of a carrier by numerically integrating the acceleration and angular velocity outputs of IMUs. However, these methods are highly sensitive to measurement accuracy, as even minor noise and bias in IMU measurements accumulate over time during integration, leading to significant drift errors. To mitigate such errors, various physical priors have been introduced as corrective constraints. For instance, Pedestrian Dead Reckoning (PDR) assumes a stable gait cycle, while Zero Velocity Update (ZUPT) exploits the prior knowledge that the velocity of a foot-mounted IMU is zero during stance phases. Furthermore, multi-sensor fusion techniques have been widely adopted to enhance system robustness by incorporating auxiliary information such as gait phase detection, vision, LiDAR, or GPS signals. In these frameworks, IMU data are typically fused with external measurements through Extended Kalman Filters (EKF), Invariant Extended Kalman Filters (IEKF), or factor graph optimization for joint state estimation. Nevertheless, physics-based methods often rely on strong assumptions about motion patterns or require additional hardware support, which limits their generalization capability and deployment flexibility in complex real-world environments.
Learning-Based Methods: With the rapid development of deep learning, researchers have explored end-to-end models that learn motion mappings from large-scale IMU datasets to improve the accuracy, robustness, and cross-scenario adaptability of IO. Early representative works such as RIDI[RIDI] and PDRNet[PDRNet] first classify the mounting position of IMU and then regress the human’s velocity or displacement. Subsequently, RoNIN identified the representation discontinuity inherent in the body-fixed frame used by RIDI—caused by changes in device orientation—and proposed the HACF, such as the global coordinate frame. RoNIN also systematically compared multiple network architectures, including TCN, LSTM, and ResNet. Following this work, most studies adopted global coordinate frame for IMU preprocessing, employing model architectures based on CNNs, LSTMs, Transformers, or their hybrid variants. For example, TLIO[TLIO] and LIDR[LIDR] extended RoNIN by integrating Stochastic Cloning Extended Kalman Filters (SCEKF) and Left Invariant Extended Kalman Filters (LIEKF), respectively; IMUNet[IMUNet], DeepILS[DeepILS], and DeepLite[Deep-Learning-for-Inertial-Positioning] focused on lightweight network design; CTIN[CTIN], SBIPTVT[SBIPTVT], and iMOT[iMOT] introduced attention mechanisms to enhance temporal modeling; RNIN-VIO, SCHNN, and SSHNN constructed multi-branch hybrid networks combining CNNs, LSTMs, and attention modules.
III Methodology
To examine the influence of different coordinate frames on pedestrian IO and to explore potential improvements in localization accuracy, this section first presents formal definitions of the body frame and the global frame in Section III-A. Section III-B then provides qualitative evidence from three complementary analyses, demonstrating that—unlike findings reported in drone-based applications—the global frame yields more discriminative representations of pedestrian motion. Building on this observation, Section III-C introduces MambaIO, which operates in the global coordinate frame and decomposes IMU measurements using a Laplacian pyramid. Dedicated modules are subsequently designed to model the implicit information embedded in each frequency component, thereby substantially enhancing the localization accuracy of pedestrian IO.
Figure 1: Schematic diagram of the coordinate system transformation relationship between the IMU and the carrier’s center of mass in UAV and pedestrian motion scenarios.
III-A Coordinate Frame Definitions
III-A1 Body Coordinate Frame
As shown in Fig.3, the IMU directly measures angular velocity and acceleration in its own body frame. These measurements serve as input data during training, while the corresponding labels (e.g., human velocity) are obtained from more accurate algorithms or external sensors (e.g., RoNIN mounted an additional smartphone on the chest to collect human velocity). For simplicity, sensor bias and noise are temporarily neglected. At time $t$, the measured quantities in the body frame are related to the true physical values as follows:
Input Data: $\left\{\begin{aligned} \boldsymbol{\omega}_{t,\text{meas}}^{\text{body}}&=\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}\\ \mathbf{a}_{t,\text{meas}}^{\text{body}}&=\mathbf{a}_{t,\text{true}}^{\text{body}}+\mathbf{R}_{t,\text{global}\to\text{body}}\,\mathbf{g}^{\text{global}}\end{aligned}\right$. (1)
Label: $\left\{\begin{aligned} \mathbf{v}_{t,\text{meas}}^{\text{body}}=\mathbf{R}_{t,\text{global}\to\text{body}}\,\mathbf{v}_{t,\text{true}}^{\text{global}}\end{aligned}\right$.
where $\boldsymbol{\omega}_{t,\text{meas}}^{\text{body}}$ and $\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}$ denote the measured and true angular velocities expressed in the body frame, respectively; $\mathbf{a}_{t,\text{meas}}^{\text{body}}$ and $\mathbf{a}_{t,\text{true}}^{\text{body}}$ denote the measured and true accelerations expressed in the body frame, respectively; $\mathbf{R}_{t,\text{global}\to\text{body}}$ is the rotation matrix that transforms vectors from the global frame to the body frame; and $\mathbf{g}^{\text{global}}=[0,0,g]^{\top}$ represents the gravitational acceleration vector in the global frame, with $g\approx 9.81\,\text{m/s}^{2}$. $\mathbf{v}_{t,\text{meas}}^{\text{body}}$ denotes the measured velocity of the carrier (e.g., human or UAV) expressed in the body frame, and $\mathbf{v}_{t,\text{true}}^{\text{global}}$ denotes the true velocity of the carrier expressed in the global frame.
III-A2 Global Coordinate Frame
The input data and labels in the global coordinate frame can be obtained by applying rotational transformations to the corresponding quantities in the body frame. Consequently, at time $t$, the model’s input data and labels are expressed as:
Input Data: $\left\{\begin{aligned} \boldsymbol{\omega}_{t,\text{meas}}^{\text{global}}&=\mathbf{R}_{t,\text{body}\to\text{global}}\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}\\ \mathbf{a}_{t,\text{meas}}^{\text{global}}&=\mathbf{R}_{t,\text{body}\to\text{global}}\mathbf{a}_{t,\text{true}}^{\text{body}}+\mathbf{g}^{\text{global}}\end{aligned}\right$. (2)
Label: $\left\{\begin{aligned} \mathbf{v}_{t,\text{meas}}^{\text{global}}=\mathbf{v}_{t,\text{true}}^{\text{global}}\end{aligned}\right$.
where $\boldsymbol{\omega}_{t,\text{meas}}^{\text{global}}$ and $\mathbf{a}_{t,\text{meas}}^{\text{global}}$ denote the angular velocity and acceleration measurements in the global frame, respectively; $\mathbf{R}_{t,\text{body}\to\text{global}}$ is the rotation matrix transforming vectors from the body frame to the global frame, typically estimated from IMU data via preintegration or pose estimation algorithms, and approximately satisfies the orthogonality condition:
$\mathbf{R}_{t,\text{body}\to\text{global}}\,\mathbf{R}_{t,\text{global}\to\text{body}}=\mathbf{I}$ (3)
where $\mathbf{I}$ denotes the identity matrix.
Fundamentally, Equations (1) and (2) are mathematically equivalent, representing the same IMU and velocity measurements in different reference frames. However, as illustrated in Fig. 3, the IMU is typically not co-located with the center of mass of the carrier (e.g., a pedestrian or a drone), causing measurements expressed in the IMU body frame to inadequately capture the true motion dynamics of the carrier.
In drone-based scenarios, the IMU is rigidly attached to the carrier, and measurements can be accurately transformed to the carrier’s center of mass using a pre-calibrated fixed transformation matrix that encodes the relative translation and rotation between the IMU and the carrier. This process is straightforward and highly accurate; consequently, some studies argue that the linear coupling between IMU measurements and pose information in Equation (1) provides a sufficient learning signal.
In contrast, for Pedestrian IO, the transformation from the IMU to the human body’s center of mass varies significantly over time and is generally infeasible to obtain in real time. For example, in RoNIN, human velocity is estimated from a smartphone mounted on the chest, while IMU measurements may originate from another phone held arbitrarily by the user. Under such practical conditions, models often directly estimate the carrier’s true velocity from raw IMU readings, implicitly learning the complex, time-varying mapping between the IMU location and an approximate human center of mass.
In this context, the linear coupling described in Equation (1) is insufficient to fully characterize the non-rigid, time-varying nature of this mapping. By comparison, the nonlinear coupling embodied in Equation (2) offers a more appropriate representation of human motion. Moreover, the global coordinate frame explicitly aligns measurements to a consistent reference orientation, effectively mitigating disturbances caused by rapid changes in carrier heading and thereby yielding motion representations that are more consistent and semantically meaningful. This observation aligns with the findings reported in RoNIN, which attributes severe discontinuities in motion representation to the drastic frame-to-frame orientation variations inherent in the body-fixed frame.
In summary, under clearly defined coordinate systems and data collection protocols, using raw IMU measurements in the body frame to estimate carrier velocity is viable for drone IO. However, for pedestrian IO, the transformation between the IMU and the human center of mass is highly dynamic and difficult to model, leading to significant representational discontinuities—a phenomenon consistent with prior empirical observations. Fig. 4 further illustrates the fundamental differences between these two IO learning paradigms.
III-B Representation Analysis
To systematically evaluate the impact of different coordinate frames on the feature representational capacity in pedestrian inertial odometry (IO) and to validate the analysis presented in Section III-A, we conduct comparative experiments using Principal Component Analysis (PCA), t-SNE, and trajectory inference visualization. All experiments are performed on six publicly available pedestrian inertial datasets. Key findings are summarized below; a complete analysis is provided in the appendix.
Figure 2: Comparison of metrics on the xxx dataset. Scatter points closer to the bottom-left corner indicate the algorithm has lower ATE and RTE, i.e., higher localization accuracy.
PCA constructs orthogonal principal components to replace correlated variables in high-dimensional data, thereby enabling dimensionality reduction and interpretable structure analysis. The analysis pipeline for inertial data under different coordinate frames is as follows: (1) concatenate the latent features extracted by the RoNIN-ResNet model from all test samples into a high-dimensional feature matrix; (2) apply PCA to this matrix and compute normalized singular values, each corresponding to the variance explained by one principal component; (3) calculate the cumulative explained variance (i.e., energy coverage ratio) of the top $k=50$ principal components to quantify the efficiency of capturing essential motion information. As shown in Fig. 3, under the RoNIN-ResNet model, the global coordinate frame (red curve) consistently achieves higher cumulative explained variance than the body-fixed frame (black curve), attaining greater energy coverage with the same number of principal components. This indicates that the global frame concentrates critical motion information more effectively—a trend consistently observed across all other datasets (see Appendix).
Figure 3: Comparison of metrics on the xxx dataset. Scatter points closer to the bottom-left corner indicate the algorithm has lower ATE and RTE, i.e., higher localization accuracy.
t-SNE maps high-dimensional features into a two- or three-dimensional space while preserving local neighborhood relationships, making it suitable for qualitatively evaluating feature separability. We extract features from raw IMU data in both the body-fixed and global frames on the RoNIN dataset and apply t-SNE for dimensionality reduction, coloring each sample according to its ground-truth speed magnitude. As shown in Fig. 4, features in the body-fixed frame exhibit substantial mixing across different speed levels with indistinct boundaries, suggesting a limited ability to encode speed-related motion semantics. In contrast, features in the global frame naturally cluster by speed, forming well-separated, coherent regions that enhance feature separability and provide a more discriminative representation for downstream models. Consistent patterns are observed across all other datasets (see Appendix).
To further validate the influence of coordinate frame selection on practical localization performance, we train and evaluate the RoNIN-ResNet model using data represented in both coordinate frames, and compare the alignment between predicted and ground-truth trajectories. As shown in Fig. 5, trajectories inferred from the global frame exhibit superior smoothness and positional accuracy compared with those from the body-fixed frame, more faithfully reconstructing the true motion state. This observation is highly consistent with the preceding PCA and t-SNE analyses, demonstrating that enhanced feature representation directly translates into improved motion modeling capability. Additional trajectory visualizations across different network architectures (e.g., CNN, Transformer) and datasets are provided in the Appendix, all of which further corroborate this conclusion.
Synthesizing evidence from the three experimental perspectives, we derive the following key conclusions: the assumption of “linear superposition of motion information and pose variation” inherent in the body-fixed frame fails to capture the complex, non-rigid, and nonlinear relative motion between pedestrians and IMUs, resulting in structurally deficient feature representations. In contrast, the global coordinate frame substantially enhances both feature compactness and class separability, producing a more discriminative and consistent latent representation for pedestrian IO. This representational advantage ultimately leads to more accurate and robust trajectory estimation, thereby validating both the effectiveness and necessity of adopting the global coordinate frame in pedestrian inertial odometry.
III-C MambaIO
Motivated by the analyses in the preceding sections, which indicate that the global coordinate frame provides superior feature representations for pedestrian IO, we further investigate strategies to enhance localization accuracy within this frame. We observe that end-to-end learning directly on raw IMU measurements is susceptible to interference from mixed signal components, including high-frequency noise and redundant local motions, which impedes the model’s ability to focus on salient motion cues. To address this, we propose MambaIO: a novel architecture that first decomposes IMU signals in the global frame into multi-scale frequency components using a Laplacian pyramid, separating the low-frequency component, which primarily encodes global motion trends, from the high-frequency component, which captures local dynamic details. Subsequently, a multi-path convolutional module, specialized for local modeling, processes the high-frequency component, while a Mamba module, adept at long-range sequential modeling, handles the low-frequency component. The outputs from both pathways are then fused to enable precise motion state inference. The overall architecture of MambaIO draws inspiration from the design principles of ConvNeXt, and its schematic diagram is presented in Fig. 6.
Figure 4: Schematic diagram of the proposed MambaIO architecture, including the Laplacian pyramid and the multi-path convolutional kernel Mamba module.
III-C1 Laplacian Pyramid Decomposition Module
Existing studies suggest that convolutional operations are inherently effective for capturing local spatiotemporal features, corresponding to high-frequency signal components, whereas attention mechanisms or state space models (e.g., Mamba) excel at modeling long-range dependencies associated with low-frequency trends. Therefore, explicitly decomposing IMU signals by frequency and designing dedicated modeling pathways for different bands can substantially improve the specificity and efficiency of feature learning.
The classical Laplacian pyramid generates multi-resolution low-frequency subbands through successive low-pass filtering and downsampling, followed by upsampling and residual computation to extract high-frequency details at each level. However, directly integrating external frequency decomposition tools, such as FFT-based filters, into deep learning frameworks often results in non-differentiable operations, training instability, or inefficient inference. To circumvent these issues while preserving the essence of multi-scale decomposition, we design a lightweight, differentiable approximation using native PyTorch operations.
Specifically, given the input feature $X\in\mathbb{R}^{N\times L}$ in the global coordinate frame, where $N=6$ denotes the three-axis acceleration and three-axis angular velocity, and $L=200$ is the number of samples in the temporal window, we first employ depthwise convolution to approximate low-pass filtering. The convolution kernel weights are fixed to a normalized averaging kernel $\mathbf{K}_{\text{avg}}\in\mathbb{R}^{k}$, satisfying $\sum_{i}\mathbf{K}_{\text{avg}}(i)=1$, and a stride $s=2$ is applied to compress the sequence length, thereby jointly approximating low-pass filtering and downsampling. In our experiments, we set the kernel size $k=5$, yielding the downsampled low-frequency feature:
$X_{ld}=\text{DWConv}_{\text{avg}}(X)\in\mathbb{R}^{N\times\lfloor L/s\rfloor}$ (4)
This design enables differentiable and efficient low-pass filtering with negligible computational overhead, avoiding gradient truncation. Next, we apply nearest-neighbor upsampling to $X_{ld}$ to restore the original sequence length, obtaining a smooth low-frequency component:
$X_{\text{low}}=\text{NearestUpSample}(X_{ld})\in\mathbb{R}^{N\times L}.$ (5)
Here, $\text{NearestUpSample}(\cdot)$ denotes the nearest-neighbor upsampling operation. Finally, the high-frequency detail component is obtained by computing the residual between the original input and the low-frequency component:
$X_{\text{high}}=X-X_{\text{low}}\in\mathbb{R}^{N\times L}$ (6)
This decomposition pipeline is fully differentiable, computationally efficient, and effectively decouples local dynamics from global trends in IMU signals, providing a robust foundation for subsequent pathway-specific modeling.
III-C2 Multi-Path Convolution Module
The Laplacian pyramid isolates the high-frequency component $X_{\text{high}}$, which primarily encodes local motion details. To efficiently extract local spatiotemporal features from this component, we design a lightweight Multi-Path Convolution (MPC) module. This module employs three parallel depthwise convolutions with distinct kernel sizes ($k_{1}=1$, $k_{2}=3$, $k_{3}=7$) to capture motion patterns across multi-scale local receptive fields. The outputs from all paths are concatenated along the channel dimension to form a comprehensive local representation:
$X_{0}$ $=\text{DWConv}_{k=1}(X_{\text{high}})\in\mathbb{R}^{N\times L}$ (7)
$X_{1}$ $=\text{DWConv}_{k=3}(X_{\text{high}})\in\mathbb{R}^{N\times L}$
$X_{2}$ $=\text{DWConv}_{k=7}(X_{\text{high}})\in\mathbb{R}^{N\times L}$
$X_{\text{concat}}$ $=\text{Concat}(X_{0},X_{1},X_{2})\in\mathbb{R}^{3N\times L}$
where DWConv denotes depthwise convolution with learnable weights, $X_{i}\in\mathbb{R}^{N\times L}$ is the output of the $i$-th path, and $\text{Concat}(\cdot)$ represents concatenation along the channel dimension.
Following the design principles of RepViT, we enhance cross-channel interaction by inserting a lightweight Squeeze-and-Excitation (SE) block after concatenation, followed by a pointwise convolution for channel fusion, yielding the final output of the high-frequency branch:
$X_{\text{MPC}}=\text{Conv}\bigl(\text{SEBlock}(X_{\text{concat}})\bigr)\in\mathbb{R}^{N\times L}$ (8)
where $X_{\text{MPC}}$ denotes the output of the MPC module, serving as the enriched local feature representation for subsequent fusion with the low-frequency pathway.
III-C3 Mamba Module
State Space Models (SSM), inspired by continuous dynamical systems, have been widely adopted to model long-range dependencies in sequential data, achieving strong performance in computer vision and sequential modeling tasks. Mamba extends the classical SSM by incorporating an input-dependent selective mechanism, enabling linear-complexity modeling of global context while effectively focusing on salient information.
State Space Model: The standard continuous-time SSM is defined as:
$\dot{h}(t)$ $=Ah(t)+Bx(t)$ (9)
$y(t)$ $=Ch(t)$ (10)
where $x(t)\in\mathbb{R}^{S\times D}$ denotes the input sequence (with $S$ as the sequence length and $D$ as the number of channels), $h(t)\in\mathbb{R}^{H}$ is the hidden state of size $H$, $y(t)\in\mathbb{R}^{S\times D}$ is the output, and $A\in\mathbb{R}^{H\times H}$, $B\in\mathbb{R}^{H\times D}$, $C\in\mathbb{R}^{D\times H}$ are learnable parameters.
Discretization: For discrete sequence processing, the system is typically discretized using Zero-Order Hold. Given a sampling interval $\Delta$, the discretized parameters are:
$\tilde{A}$ $=\exp(\Delta A)$ (11)
$\tilde{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot(\Delta B)$ (12)
yielding the recurrence:
$h_{t}$ $=\tilde{A}h_{t-1}+\tilde{B}x_{t}$ (13)
$y_{t}$ $=Ch_{t}$ (14)
[…]
Mamba enhances this framework with an input-dependent selection mechanism that improves its ability to attend to relevant contextual cues. As the internal mechanics of Mamba are not central to this work, we omit further details. We implement the Mamba module following the MambaVision construction, as illustrated in Fig. 7. Based on the State Space Discretization (SSD) procedure, the computation in the Mamba module can be abstracted as:
$X_{1}$ $=\text{SSD}\bigl(\sigma(\text{Conv}(\text{Linear}(X_{\text{low}})))\bigr)$ (15)
$X_{2}$ $=\sigma(\text{Conv}(\text{Linear}(X_{\text{low}})))$ (16)
$X_{\text{Mamba}}$ $=\text{Linear}\bigl(\text{Concat}(X_{1},X_{2})\bigr)$ (17)
where $\sigma$ denotes a nonlinear activation function and Linear represents a linear projection layer.
This module captures global motion dependencies from the low-frequency component $X_{\text{low}}$ with $\mathcal{O}(L)$ linear complexity, complementing the high-frequency MPC module. Prior work indicates that adding attention mechanisms in deeper layers can further improve modeling accuracy; therefore, a standard attention module is appended after the final Mamba block to enhance the representational weight of critical frames.
Finally, the outputs of the MPC module ($X_{\text{MPC}}$) and the Mamba module ($X_{\text{Mamba}}$) are fused to obtain the final representation of the dual-path Laplacian pyramid architecture, which is subsequently used for pose regression:
$X_{\text{out}}=\text{Conv}(\text{Concat}(X_{\text{MPC}},X_{\text{Mamba}}))\in\mathbb{R}^{N\times L}$ (18)
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
We conduct experiments on six publicly available pedestrian inertial odometry datasets: RIDI, RoNIN, RNIN, OxIOD, TLIO, and IMUNet. These datasets cover diverse indoor and outdoor environments and include a wide range of motion patterns, such as walking, running, and stair climbing, providing a comprehensive basis for evaluating model generalization and robustness.
IV-A2 Baseline Methods
Considering that learning-based approaches substantially outperform traditional physics-driven methods in localization accuracy, we select representative deep learning models as baselines. These include CNN-based methods: RoNIN-ResNet, IMUNet, DeepLite, DeepILS, and the network proposed in TLIO; attention- or LSTM-based methods: RoNIN-LSTM, CTIN, iMOT, and SPBIT; and hybrid architectures, such as the network proposed in RNIN. In addition, following the practices in IMUNet and DeepLite, we include general-purpose networks that have demonstrated strong performance in other domains, including RepViT, MobileNetv3, and MobileNetV4, to enable fair cross-architecture comparison.
IV-A3 Evaluation Metrics
We adopt two widely used metrics in inertial odometry: Absolute Trajectory Error (ATE) and Relative Trajectory Error (RTE). ATE computes the root-mean-square error between the estimated trajectory and the ground-truth trajectory after global alignment, reflecting overall localization accuracy. RTE measures trajectory error within a fixed-length sliding window, evaluating short-term consistency and drift control.
IV-A4 Training Details
The model employs four hierarchical feature extraction stages with channel dimensions [64, 128, 256, 512]. We optimize the network using Adam with an initial learning rate of $10^{-4}$ and terminate training early if the learning rate decays to $10^{-6}$. The maximum number of training epochs is set to 40, based on validation set observations indicating that most models begin to overfit beyond this point. All experiments are conducted on five NVIDIA RTX 3090 GPUs using CUDA 12.2 and PyTorch 2.5.1.
IV-B Coordinate Frame Representation Comparison
Figure 8 presents the ATE and RTE results of all baseline methods when trained and evaluated under the body-fixed and global coordinate frames separately. It is evident that, in the body-fixed frame, most models struggle to learn effective motion representations, resulting in substantially lower localization accuracy compared to their performance in the global frame. This finding is consistent with the trajectory visualization analysis in  III-B, which shows that the global coordinate frame provides a more stable and discriminative motion representation, thereby markedly enhancing model performance. Detailed comparative results across additional datasets are provided in the Appendix.
IV-C Algorithm Comparison
Table 1 summarizes the quantitative results of all methods on the six public pedestrian IO datasets. The experiments demonstrate that the proposed MambaIO consistently outperforms existing state-of-the-art approaches across most scenarios. For example, on the RoNIN dataset, MambaIO reduces ATE by approximately x% and RTE by about x% compared to RoNIN-ResNet. On the largest dataset, TLIO, MambaIO further improves over the latest Transformer-based inertial localization method, reducing ATE by x%–x% and RTE by x%–x%. These results quantitatively confirm the significant gains in localization accuracy achieved by our method.
To provide a visual assessment of localization performance, we present representative trajectories from each dataset (Fig. 9). For simple straight-line or gently curved trajectories (Fig. 9(a)–(c)), both RoNIN-ResNet and MambaIO closely follow the ground-truth paths. In contrast, for trajectories with multiple turns or complex dynamics (Fig. 9(d)–(f)), RoNIN-ResNet exhibits noticeable deviations, whereas MambaIO maintains high accuracy. This improvement is attributed to the Laplacian pyramid’s multi-scale decomposition of IMU signals, which allows the model to jointly capture local details and global trends, resulting in a more robust and precise motion representation.
IV-D Ablation Study
To evaluate the effectiveness of the proposed components, we construct two ablation variants: (1) Conv-only, which processes raw IMU signals using only the multi-path convolution module; and (2) Mamba-only, which processes raw IMU signals using only the Mamba module. Table 2 presents the performance of these variants, the full MambaIO model, and RoNIN-ResNet across the six datasets.
The results indicate that while Conv-only achieves basic localization capability, its accuracy is substantially lower than that of the full MambaIO, suggesting that local modeling alone is insufficient to capture the complex multi-scale dynamics in IMU signals. Similarly, Mamba-only underperforms, highlighting the limited ability of purely global modeling to represent high-frequency local details such as gait transients or turning jitter. Notably, both ablation variants still outperform RoNIN-ResNet, confirming the intrinsic effectiveness of the designed modules. The full MambaIO achieves the best performance across all datasets, significantly surpassing current state-of-the-art methods, which strongly validates the complementary benefits of the high/low-frequency decomposition strategy and the dual-path collaborative architecture.
In summary, the ablation study not only confirms the necessity of each component but also underscores the critical role of multi-scale feature disentanglement and joint local–global modeling in enhancing pedestrian inertial localization accuracy.
V Conclusion
This work systematically investigates the influence of coordinate frame selection on pedestrian inertial odometry, highlighting the representational advantages of the global coordinate frame in non-rigid carrier scenarios. Building on this insight, we propose MambaIO, which employs a Laplacian pyramid to separate high- and low-frequency components of IMU signals and integrates a multi-path convolution module with a Mamba module to jointly model local details and global motion trends. Despite Mamba’s linear complexity, which is lower than that of conventional Transformers, its inference speed remains relatively limited when processing long sequences, particularly in resource-constrained mobile deployment settings. Future research will focus on developing even lighter and more efficient sequence modeling strategies to further improve real-time performance while maintaining accuracy.