### v2

# MambaIO: Global-Coordinate Inertial Odometry for Pedestrians via Multi-Scale Frequency-Decoupled Modeling

Abstract
Inertial Odometry (IO) enables real-time localization using only acceleration and angular velocity measurements from an Inertial Measurement Unit (IMU), making it a promising solution for localization in consumer-grade applications. Traditionally, researchers have routinely transformed IMU measurements into the global frame to obtain smoother motion representations. However, recent studies in drone scenarios have demonstrated that the body frame can significantly improve localization accuracy, prompting a re-evaluation of the suitability of the global frame for pedestrian IO. To address this issue, this paper systematically evaluates the effectiveness of the global frame in pedestrian IO through theoretical analysis, qualitative inspection, and quantitative experiments. Building upon these findings, we further propose MambaIO, which decomposes IMU measurements into high-frequency and low-frequency components using a Laplacian pyramid. The low-frequency component is processed by a Mamba architecture to extract implicit contextual motion cues, while the high-frequency component is handled by a convolutional structure to capture fine-grained local motion details. Experiments on multiple public datasets show that MambaIO substantially reduces localization error and achieves state-of-the-art (SOTA) performance. To the best of our knowledge, this is the first application of the Mamba architecture to the IO task.
I Introduction
The Inertial Measurement Unit (IMU) serves as the primary data source for Inertial Odometry (IO), providing acceleration and angular velocity measurements[Tartan-IMU]. Owing to its low cost, compact size, and minimal power consumption, the IMU is widely integrated into mobile devices such as smartphones, robots, and unmanned aerial vehicles (UAV)[AirIO].
IO aims to estimate the trajectory of a moving carrier (e.g., pedestrian or UAV) based solely on acceleration and angular velocity measurements from the IMU, enabling high-precision self-localization and tracking[surveyILS]. Traditional IO approaches primarily rely on numerical integration, but noise and bias inherent in IMU measurements accumulate over time, leading to significant drift errors[SINS]. In contrast, learning-based IO methods, trained end-to-end on large-scale IMU datasets, automatically uncover latent motion regularities and demonstrate superior performance in terms of localization accuracy and robustness, and cross-scenario generalizability[RoNIN].
Figure 1: Performance comparison of algorithms on the RoNIN dataset in terms of ATE and RTE. Points closer to the lower-left corner indicate lower errors and thus higher localization accuracy.
Early learning-based IO methods attempted to estimate the trajectory of the carrier directly in the body frame[6-DOF-IO]. However, RoNIN reported that the body frame undergoes frequent changes during motion, resulting in discontinuous motion representations and associated issues. To mitigate these problems, RoNIN introduced a Heading-Agnostic Coordinate Frame (HACF), defined as any coordinate system whose $Z$-axis is aligned with the gravity direction, thereby providing a more stable and reliable motion representation[RoNIN]. This approach has since become the prevailing data processing paradigm in the IO community. Nevertheless, recent studies in drone-based scenarios have shown that, under certain conditions, using the body frame directly can lead to substantially improved localization accuracy[AirIO]. This counterintuitive observation motivates a reassessment of a fundamental question:
Is the global frame representation truly beneficial for pedestrian IO?
To address this question, we first conduct a systematic analysis of the representational characteristics of inertial data under different coordinate frames. Through kinematic analysis and visualization, we verify that the global frame provides more discriminative feature representations for pedestrian motion. Building on this finding, we propose MambaIO, which decomposes IMU measurements into frequency components to enable targeted learning. Experimental results demonstrate that MambaIO further improves localization accuracy in pedestrian IO, with selected results shown in Fig. 1. The main contributions of this work are summarized as follows:
  • Coordinate Frame Representation Analysis: We systematically analyze the differences in motion representation across coordinate frames, validating the necessity of adopting the global frame in pedestrian IO.
  • Laplacian Pyramid Decomposition: We propose a PyTorch-based Laplacian pyramid method that efficiently decomposes IMU measurements into high- and low-frequency components, thereby enhancing the specificity and effectiveness of subsequent feature learning.
  • Mamba Module: We introduce the Mamba architecture to specifically model the contextual motion information embedded in the low-frequency component.
  • Multi-Scale Convolutional Module: We design a multi-scale convolutional module to extract local motion features from the high-frequency component, which collaborates with the Mamba module to achieve comprehensive modeling of motion dynamics.
II Related Work
From a methodological perspective, research on IO can be broadly divided into two categories: traditional methods based on Newtonian mechanics and data-driven approaches leveraging machine learning[SurveyofIndoorInertial].
Newtonian Mechanics-Based Methods: Early approaches, exemplified by Strapdown Inertial Navigation Systems (SINS), estimate the position and orientation of a carrier by numerically integrating the acceleration and angular velocity outputs of IMUs[SINS]. However, these methods are highly sensitive to measurement accuracy, as even minor noise and bias in IMU measurements accumulate over time during integration, leading to significant drift errors[EqNIO]. To mitigate such errors, various physical priors have been introduced as corrective constraints. For instance, Pedestrian Dead Reckoning (PDR) assumes a stable gait cycle[PDR], while Zero Velocity Update (ZUPT) exploits the prior knowledge that the velocity of a foot-mounted IMU is zero during stance phases[AdaptiveThreshold-BasedZUPT]. Furthermore, multi-sensor fusion techniques have been widely adopted to enhance system robustness by incorporating auxiliary information such as gait phase detection[PDR], vision[A-visual-inertial-approach-to-human-gait-estimation], or LiDAR[FAST-LIVO2]. In these frameworks, IMU data are typically fused with external measurements through Extended Kalman Filters (EKF)[RNIN-VIO, MSCKF] for joint state estimation. Nevertheless, physics-based methods often rely on strong assumptions about motion patterns or require additional hardware support, which limits their generalization capability and deployment flexibility in complex real-world environments[RoNIN].
Learning-Based Methods: With the rapid development of deep learning, researchers have explored end-to-end models that learn motion mappings from large-scale IMU datasets to improve the accuracy, robustness, and cross-scenario adaptability of IO[surveyILS]. Early representative works such as RIDI[RIDI] and PDRNet[PDRNet] first classify the mounting position of IMU and then regress the pedestrian’s velocity or displacement. Subsequently, RoNIN identified the representation discontinuity inherent in the body frame used by RIDI—caused by changes in device orientation—and proposed the HACF, such as the global frame[RoNIN]. RoNIN also systematically compared multiple network architectures, including TCN, LSTM, and ResNet. Following this work, most studies adopted global frame for IMU preprocessing, employing model architectures based on CNNs, LSTMs, Transformers, or their hybrid variants. For example, TLIO[TLIO] and LIDR[LIDR] extended RoNIN by integrating Stochastic Cloning Extended Kalman Filters (SCEKF) and Left Invariant Extended Kalman Filters (LIEKF), respectively; IMUNet[IMUNet], DeepILS[DeepILS], and DeepLite[Deep-Learning-for-Inertial-Positioning] focused on lightweight network design; CTIN[CTIN], SBIPTVT[SBIPTVT], and iMOT[iMOT] introduced attention mechanisms to enhance temporal modeling; RNIN-VIO[RNIN-VIO], SCHNN[SCHNN], and SSHNN[SSHNN] constructed multi-branch hybrid networks combining CNNs, LSTMs, and attention modules.
Figure 2: Schematic diagram of the coordinate system transformation relationship between the IMU and the carrier’s center of mass in UAV and pedestrian motion scenarios.
III Methodology
This section first provides formal definitions of IMU measurements and their corresponding labels in learning-based IO. Subsequently, through kinematic analysis and visualization-based comparison, we validate the necessity of the global frame in pedestrian IO—a conclusion that contrasts with prior findings[AirIO] in drone-based scenarios. Building on this insight, we propose MambaIO, a novel framework designed to more accurately learn the implicit motion states information embedded in IMU measurements within the global frame, thereby further enhancing pedestrian IO performance.
III-A Coordinate Frame Definitions
III-A1 body frame
As shown in Fig. 2, the IMU directly measures angular velocity and acceleration in its own body frame. These measurements serve as training data, while the corresponding labels (e.g., pedestrian or UAV velocity) are obtained from more accurate algorithms or external sensors (e.g., RoNIN mounted an additional smartphone on the chest to collect pedestrian velocity[RoNIN]). For simplicity, sensor bias and noise are temporarily neglected. At time $t$, the measured quantities in the body frame are related to the true physical values as follows:
Training Data: $\left\{\begin{aligned} \boldsymbol{\omega}_{t,\text{meas}}^{\text{body}}&=\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}\\ \mathbf{a}_{t,\text{meas}}^{\text{body}}&=\mathbf{a}_{t,\text{true}}^{\text{body}}+\mathbf{R}_{t,\text{global}\to\text{body}}\,\mathbf{g}^{\text{global}}\end{aligned}\right$. (1)
Label: $\left\{\begin{aligned} \mathbf{v}_{t,\text{meas}}^{\text{body}}=\mathbf{R}_{t,\text{global}\to\text{body}}\,\mathbf{v}_{t,\text{true}}^{\text{global}}\end{aligned}\right$.
where $\boldsymbol{\omega}_{t,\text{meas}}^{\text{body}}$ and $\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}$ denote the measured and true angular velocities expressed in the body frame, respectively; $\mathbf{a}_{t,\text{meas}}^{\text{body}}$ and $\mathbf{a}_{t,\text{true}}^{\text{body}}$ denote the measured and true accelerations expressed in the body frame, respectively; $\mathbf{R}_{t,\text{global}\to\text{body}}$ is the rotation matrix that transforms vectors from the global frame to the body frame; and $\mathbf{g}^{\text{global}}=[0,0,9.81]^{\top}$ represents the gravitational acceleration vector in the global frame. $\mathbf{v}_{t,\text{meas}}^{\text{body}}$ denotes the measured velocity of the carrier (e.g., pedestrian or UAV) expressed in the body frame, and $\mathbf{v}_{t,\text{true}}^{\text{global}}$ denotes the true velocity of the carrier expressed in the global frame.
Figure 3: Comparison of IO Learning Processes in body frame and global frame.
III-A2 global frame
The training data and labels in the global frame can be obtained by applying rotational transformations to the corresponding quantities in the body frame. Consequently, at time $t$, the model’s training data and labels are expressed as:
Training Data: $\left\{\begin{aligned} \boldsymbol{\omega}_{t,\text{meas}}^{\text{global}}&=\mathbf{R}_{t,\text{body}\to\text{global}}\boldsymbol{\omega}_{t,\text{true}}^{\text{body}}\\ \mathbf{a}_{t,\text{meas}}^{\text{global}}&=\mathbf{R}_{t,\text{body}\to\text{global}}\mathbf{a}_{t,\text{true}}^{\text{body}}+\mathbf{g}^{\text{global}}\end{aligned}\right$. (2)
Label: $\left\{\begin{aligned} \mathbf{v}_{t,\text{meas}}^{\text{global}}=\mathbf{v}_{t,\text{true}}^{\text{global}}\end{aligned}\right$.
where $\boldsymbol{\omega}_{t,\text{meas}}^{\text{global}}$ and $\mathbf{a}_{t,\text{meas}}^{\text{global}}$ denote the angular velocity and acceleration measurements in the global frame, respectively; $\mathbf{R}_{t,\text{body}\to\text{global}}$ is the rotation matrix transforming vectors from the body frame to the global frame, typically estimated from IMU data via preintegration or pose estimation algorithms, and approximately satisfies the orthogonality condition:
$\mathbf{R}_{t,\text{body}\to\text{global}}\,\mathbf{R}_{t,\text{global}\to\text{body}}=\mathbf{I}$ (3)
where $\mathbf{I}$ denotes the identity matrix.
Fundamentally, Equations (1) and (2) are mathematically equivalent, representing the same IMU and velocity measurements in different reference frames. However, as illustrated in Fig. 2, the IMU is typically not co-located with the center of mass of the carrier (e.g., pedestrian or UAV), causing measurements expressed in the IMU body frame to inadequately capture the true motion dynamics of the carrier.
In drone-based scenarios, the IMU is rigidly attached to the carrier, and measurements can be accurately transformed to the carrier’s center of mass using a pre-calibrated fixed transformation matrix that encodes the relative translation and rotation between the IMU and the carrier. This process is straightforward and highly accurate; consequently, some studies argue that the linear coupling between IMU measurements and pose information in Equation (1) provides a sufficient learning signal[AirIO].
In contrast, for Pedestrian IO, the transformation from the IMU to the pedestrian body’s center of mass varies significantly over time and is generally infeasible to obtain in real time. For example, in RoNIN, pedestrian velocity is estimated from a smartphone mounted on the chest, while IMU measurements may originate from another phone held arbitrarily by the user[RoNIN]. Under such practical conditions, models often directly estimate the carrier’s true velocity from raw IMU readings, implicitly learning the complex, time-varying mapping between the IMU location and an approximate pedestrian center of mass.
In this context, the linear coupling described in Equation (1) is insufficient to fully characterize the non-rigid, time-varying nature of this mapping. By comparison, the nonlinear coupling embodied in Equation (2) offers a more appropriate representation of pedestrian motion. Moreover, the global frame explicitly aligns measurements to a consistent reference orientation, effectively mitigating disturbances caused by rapid changes in carrier heading and thereby yielding motion representations that are more consistent and semantically meaningful[RoNIN]. This observation aligns with the findings reported in RoNIN[RoNIN], which attributes severe discontinuities in motion representation to the drastic frame-to-frame orientation variations inherent in the body frame.
In summary, under clearly defined coordinate systems and data collection protocols, using raw IMU measurements in the body frame to estimate carrier velocity is viable for drone IO. However, for pedestrian IO, the transformation between the IMU and the pedestrian center of mass is highly dynamic and difficult to model, leading to significant representational discontinuities—a phenomenon consistent with prior empirical observations. Fig. 3 further illustrates the fundamental differences between these two IO learning paradigms.
III-B Representation Analysis
To systematically validate the analysis presented in Section III-A, this section employs t-SNE visualization and trajectory inference visualization to conduct a comparative analysis of motion representations under different coordinate frames.
t-SNE maps high-dimensional features into a two- or three-dimensional space while preserving local neighborhood relationships, making it suitable for qualitative assessment of feature separability. In our experiments, we apply t-SNE to IMU measurements represented in both the body and global frames, coloring each sample according to its ground-truth speed magnitude. As shown in Fig. 4, features in the body frame exhibit significant overlap across different speed ranges, with blurred boundaries, indicating limited capacity to encode speed-related motion semantics. In contrast, features in the global frame naturally cluster by speed, forming distinct, well-separated, and structurally coherent regions that substantially enhance feature discriminability and provide more informative representations for downstream models. This consistent pattern is observed across all other datasets (see Appendix for details).
Trajectory Visualization: To intuitively assess the impact of coordinate frame representations on learning in IO, we plot the trajectories inferred by the same network using identical IMU data but processed in either the body or global frame. As shown in Fig. 5, the red curve corresponding to the global frame closely aligns with the green ground-truth trajectory, whereas the blue curve representing the body frame deviates more significantly. This indicates that trajectories inferred from the global frame exhibit higher smoothness and positional accuracy, enabling more faithful reconstruction of the actual motion. This observation is highly consistent with the preceding t-SNE analyses, demonstrating that enhanced feature representation directly translates into improved motion modeling performance. This consistent pattern is observed across all other datasets (see Appendix for details).
In summary, through kinematic analysis and multi-perspective visualization comparisons, we conclude that the global frame facilitates more accurate and robust trajectory estimation, thereby validating its effectiveness and necessity in pedestrian IO.
Figure 4: t-SNE visualization of features from the RIDI dataset, demonstrating that the global frame representation yields better separability of motion patterns.
III-C MambaIO
The preceding analysis and experiments demonstrate that the global frame yields motion representations with higher observability. However, an effective method for accurately modeling the implicit pose transformations and motion information embedded in IMU measurements remains lacking. We argue that IMU signals contain multi-scale motion characteristics mixed with noise and other interfering components, which collectively hinder a model’s ability to focus on the most relevant information. To address this challenge, we propose MambaIO, which first decomposes IMU measurements in the frequency domain and then extracts informative features tailored to the distinct characteristics of each frequency band. The overall architecture is illustrated in Fig. 6.
III-C1 Laplacian Pyramid Decomposition Module
Existing studies suggest that convolutional operations are inherently effective at capturing local spatiotemporal features, which correspond to high-frequency signal components, whereas attention mechanisms or state space models (e.g., Mamba) excel at modeling long-range dependencies associated with low-frequency trends. Therefore, explicitly decomposing IMU signals by frequency and designing dedicated modeling pathways for different bands can substantially improve the specificity and efficiency of feature learning.
Figure 5: Trajectory predictions by RoNIN ResNet on the RoNIN dataset under Body (blue) and Global (orange) coordinate representations. The Global-frame trajectory aligns more closely with the ground truth (green), indicating improved localization accuracy.
Classical frequency decomposition methods, such as the Laplacian pyramid, generate multi-resolution low-frequency subbands through successive low-pass filtering and downsampling, followed by upsampling and residual computation to extract high-frequency details at each level. The visualization of this process is shown in Fig. 6(b). However, directly integrating external frequency decomposition tools—such as FFT-based filters—into deep learning frameworks often introduces non-differentiable operations, training instability, or inefficient inference. To circumvent these issues while preserving the essence of multi-scale decomposition, we design a lightweight, fully differentiable approximation using native PyTorch operations.
Specifically, given the input feature $X\in\mathbb{R}^{N\times L}$ in the global frame, where $N=6$ denotes the three-axis acceleration and three-axis angular velocity, and $L=200$ is the number of samples in the temporal window, we first employ depthwise convolution to approximate low-pass filtering. The convolution kernel weights are fixed to a normalized averaging kernel $\mathbf{K}_{\text{avg}}\in\mathbb{R}^{k}$, satisfying $\sum_{i}\mathbf{K}_{\text{avg}}(i)=1$, and a stride $s=2$ is applied to compress the sequence length, thereby jointly approximating low-pass filtering and downsampling. In our experiments, we set the kernel size $k=5$, yielding the downsampled low-frequency feature:
$X_{ld}=\text{DWConv}_{\text{avg}}(X)\in\mathbb{R}^{N\times\lfloor L/s\rfloor}.$ (4)
This design enables differentiable and efficient low-pass filtering with negligible computational overhead, avoiding gradient truncation. Next, we apply nearest-neighbor upsampling to $X_{ld}$ to restore the original sequence length, obtaining a smooth low-frequency component:
$X_{\text{low}}=\text{NearestUpSample}(X_{ld})\in\mathbb{R}^{N\times L}.$ (5)
Here, $\text{NearestUpSample}(\cdot)$ denotes the nearest-neighbor upsampling operation. Finally, the high-frequency detail component is obtained by computing the residual between the original input and the low-frequency component:
$X_{\text{high}}=X-X_{\text{low}}\in\mathbb{R}^{N\times L}.$ (6)
This decomposition pipeline is fully differentiable, computationally efficient, and effectively decouples local dynamics from global trends in IMU signals, providing a robust foundation for subsequent pathway-specific modeling.
Figure 6: Schematic diagram of the proposed MambaIO architecture, including the Laplacian pyramid and the multi-path convolutional kernel Mamba module.
III-C2 Multi-Path Convolution Module
The Laplacian pyramid isolates the high-frequency component $X_{\text{high}}$, which primarily encodes local motion details. To efficiently extract local spatiotemporal features from this component, we design a lightweight Multi-Path Convolution (MPC) module. This module employs three parallel depthwise convolutions with distinct kernel sizes ($k_{1}=1$, $k_{2}=3$, $k_{3}=7$) to capture motion patterns across multi-scale local receptive fields. The outputs from all paths are concatenated along the channel dimension to form a comprehensive local representation:
$X_{0}$ $=\text{DWConv}_{k=1}(X_{\text{high}})\in\mathbb{R}^{N\times L},$ (7)
$X_{1}$ $=\text{DWConv}_{k=3}(X_{\text{high}})\in\mathbb{R}^{N\times L},$
$X_{2}$ $=\text{DWConv}_{k=7}(X_{\text{high}})\in\mathbb{R}^{N\times L},$
$X_{\text{concat}}$ $=\text{Concat}(X_{0},X_{1},X_{2})\in\mathbb{R}^{3N\times L},$
where DWConv denotes depthwise convolution with learnable weights, $X_{i}\in\mathbb{R}^{N\times L}$ is the output of the $i$-th path, and $\text{Concat}(\cdot)$ represents concatenation along the channel dimension.
We then enhance cross-channel interaction by inserting a lightweight Squeeze-and-Excitation (SE) block after concatenation, followed by a pointwise convolution for channel fusion, yielding the final output of the high-frequency branch:
$X_{\text{MPC}}=\text{Conv}\bigl(\text{SEBlock}(X_{\text{concat}})\bigr)\in\mathbb{R}^{N\times L},$ (8)
where $X_{\text{MPC}}$ denotes the output of the MPC module, serving as the enriched local feature representation for subsequent fusion with the low-frequency pathway.
III-C3 Mamba Module
State Space Models (SSMs), inspired by continuous dynamical systems, have been widely adopted to model long-range dependencies in sequential data. Mamba extends the classical SSM by incorporating an input-dependent selective mechanism, enabling linear-complexity modeling of global context while effectively focusing on salient information[Mamba].
State Space Model: The standard continuous-time SSM is defined as:
$\dot{h}(t)$ $=Ah(t)+Bx(t),$ (9)
$y(t)$ $=Ch(t),$ (10)
where $x(t)\in\mathbb{R}^{S\times D}$ denotes the input sequence (with $S$ as the sequence length and $D$ as the number of channels), $h(t)\in\mathbb{R}^{H}$ is the hidden state of dimension $H$, $y(t)\in\mathbb{R}^{S\times D}$ is the output, and $A\in\mathbb{R}^{H\times H}$, $B\in\mathbb{R}^{H\times D}$, $C\in\mathbb{R}^{D\times H}$ are learnable parameters.
Discretization: For discrete sequences, the system is typically discretized using the Zero-Order Hold method. Given a sampling interval $\Delta$, the discretized parameters are:
$\tilde{A}$ $=\exp(\Delta A),$ (11)
$\tilde{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)(\Delta B),$ (12)
yielding the recurrence:
$h_{t}$ $=\tilde{A}h_{t-1}+\tilde{B}x_{t},$ (13)
$y_{t}$ $=Ch_{t}.$ (14)
[…]
Mamba enhances this framework with an input-dependent selection mechanism that improves its ability to attend to relevant contextual cues. Since the internal mechanics of Mamba are not central to this work, we omit further details. As illustrated in Fig. 6 (a), based on the State Space Discretization (SSD) procedure, the computation in the Mamba module can be abstracted as:
$X_{1}$ $=\text{SSD}\bigl(\sigma(\text{Conv}(\text{Linear}(X_{\text{low}})))\bigr),$ (15)
$X_{2}$ $=\sigma(\text{Conv}(\text{Linear}(X_{\text{low}}))),$ (16)
$X_{\text{Mamba}}$ $=\text{Linear}\bigl(\text{Concat}(X_{1},X_{2})\bigr),$ (17)
where $\sigma$ denotes a nonlinear activation function and Linear represents a linear projection layer.
This module captures global motion dependencies from the low-frequency component $X_{\text{low}}$ with $\mathcal{O}(L)$ linear complexity, complementing the high-frequency MPC module.
Finally, the outputs of the MPC module ($X_{\text{MPC}}$) and the Mamba module ($X_{\text{Mamba}}$) are fused to obtain the final representation of the dual-path Laplacian pyramid architecture, which is subsequently used for pose regression:
$X_{\text{out}}=\text{Conv}(\text{Concat}(X_{\text{MPC}},X_{\text{Mamba}}))\in\mathbb{R}^{N\times L}.$ (18)
TABLE I: Algorithm performance across datasets. Best and second-best results are marked in orange and yellow, respectively.
Dataset RIDI RoNIN RNIN IMUNet OxIOD Average
Metrics (Performance: meters) ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE ATE RTE
Hybrid- based MambaIO (Ours) 1.750 1.996 4.480 3.726 3.002 2.166 5.303 3.882 1.568 0.980 3.221 2.550
RNIN (ISMAR 2021)[RNIN-VIO] 2.771 2.992 6.657 4.079 4.203 2.285 8.400 5.055 3.662 4.337 5.139 3.150
CNN- based RoNIN ResNet (ICRA 2020)[RoNIN] 2.429 2.599 6.125 4.080 3.924 2.621 8.159 4.706 3.498 1.305 4.827 3.062
TLIO (RA-L 2020)[TLIO] 2.240 2.524 5.522 3.799 4.297 3.353 7.490 5.273 2.020 0.969 4.440 3.184
IMUNet (TIM 2024)[IMUNet] 3.201 3.240 6.071 4.043 3.156 2.262 8.676 5.947 1.638 1.043 4.548 3.307
DeepILS (IOT-J 2025)[DeepILS] 1.893 2.363 5.803 3.903 3.675 2.392 7.861 5.585 3.294 1.233 4.505 3.095
DeepLite (IOT-J 2025)[DeepLite] 3.658 3.814 6.363 4.382 3.878 2.892 7.459 5.746 2.048 1.112 4.681 3.589
RepVIT-M0.6 (CVPR 2024)[Repvit] 2.523 2.880 5.923 3.762 3.916 2.503 7.103 5.141 2.859 1.165 4.465 3.090
MobileNetV4-S (arXiv 2024)[MobileNetV4] 2.677 3.141 5.256 3.732 3.768 2.567 7.787 6.661 4.408 1.617 4.779 3.544
Tartan IMU (CVPR 2025)[Tartan-IMU] 2.254 2.581 5.212 3.748 3.169 1.982 6.291 4.492 1.670 0.940 3.719 2.749
Transformer- based iMOT (AAAI 2025)[iMOT] 2.788 3.485 5.967 4.856 6.359 3.229 8.953 6.616 21.802 5.197 9.174 4.677
CTIN (AAAI 2022)[CTIN] 2.325 2.918 4.878 4.063 4.732 2.813 5.589 4.522 4.846 1.563 4.474 3.176
SBIPTVTL (CSCWD 2024)[SBIPTVT] 2.074 2.317 5.300 3.799 3.781 2.269 6.466 4.325 1.732 0.954 3.871 2.733
Improvement MambaIO vs RoNIN ResNet(%) 28.0 23.2 26.9 8.7 23.5 17.4 35.0 17.5 55.2 24.9 33.3 16.7
Figure 7: Comparison of trajectory error for RoNIN ResNet under global and body coordinate frames across multiple datasets.
IV Experiments
IV-A Experimental Setup
IV-A1 Datasets
We conduct experiments on five publicly available pedestrian IO datasets: RIDI[RIDI], RoNIN[RoNIN], RNIN[RNIN-VIO], OxIOD[Oxiod], and IMUNet[IMUNet]. These datasets span diverse indoor and outdoor environments and encompass a wide range of motion patterns—including walking, running, and stair climbing—providing a comprehensive basis for evaluating model generalization and robustness.
IV-A2 Baseline Methods
Given that learning-based approaches substantially outperform traditional physics-driven methods in localization accuracy, we select representative deep learning models as baselines. These include CNN-based methods: RoNIN ResNet[RoNIN], IMUNet[IMUNet], DeepLite[DeepLite], DeepILS[DeepILS], and the network proposed in TLIO[TLIO]; attention-based methods: CTIN[CTIN], iMOT[iMOT], and SBIPTVTL[SBIPTVT]; and hybrid architectures, such as the network proposed in RNIN[RNIN-VIO] and Tartan IMU[Tartan-IMU]. In addition, following the practices in IMUNet and DeepLite, we also include general-purpose networks that have demonstrated strong performance in other domains—namely RepViT[Repvit], and MobileNetV4[MobileNetV4]—to enable fair cross-architecture comparison.
IV-A3 Evaluation Metrics
We adopt two widely used metrics in IO: Absolute Trajectory Error (ATE) and Relative Trajectory Error (RTE)[Ionet]. ATE computes the Root Mean Square Error (RMSE) between the estimated trajectory and the ground-truth trajectory after global alignment, reflecting overall localization accuracy. RTE computes the RMSE within a fixed-length sliding window, evaluating short-term consistency and drift control.
IV-A4 Training Details
The model employs four hierarchical feature extraction stages with channel dimensions [64, 128, 256, 512]. We optimize the network using the Adam optimizer with an initial learning rate of $10^{-4}$ and terminate training early if the learning rate decays to $10^{-6}$. The maximum number of training epochs is set to 40. All experiments are conducted on five NVIDIA RTX 3090 GPUs using CUDA 12.2 and PyTorch 2.5.1.
IV-B Coordinate Frame Representation Comparison
Fig. 7 presents the ATE and RTE curves of RoNIN ResNet when evaluated on IMU data represented in the body-fixed versus the global frame. It is evident that, under the body frame, RoNIN ResNet struggle to learn effective motion representations, resulting in substantially degraded localization accuracy compared to the global frame. This finding aligns with the trajectory visualization analysis in Section III-B: the global frame provides more stable and discriminative motion representations, thereby significantly enhancing model performance. Detailed comparisons across additional models are provided in the Appendix.
TABLE II: The ablation study results. The best result per metric is highlighted with orange background. All metrics are in meters.
Dataset Metric Conv-only Mamba-only MambaIO (full)
RoNIN ATE 6.835 6.063 4.480
RTE 4.160 4.000 3.726
RIDI ATE 2.318 2.114 1.750
RTE 2.656 2.419 1.996
IMUNet ATE 6.132 6.366 5.303
RTE 4.207 4.493 3.882
RNIN ATE 6.154 3.907 3.002
RTE 3.223 2.847 2.166
OxIOD ATE 1.844 2.723 1.568
RTE 1.020 1.167 0.980
IV-C Algorithm Comparison
Table I summarizes the quantitative results of all methods on the five public pedestrian IO datasets. The experiments demonstrate that the proposed MambaIO consistently outperforms existing state-of-the-art approaches across most scenarios. For example, on the RoNIN dataset, MambaIO reduces ATE by approximately $26.9\%$ and RTE by about $8.7\%$ compared to RoNIN ResNet. Furthermore, the average results across all five datasets show that MambaIO achieves the highest localization accuracy, with improvements of $33.3\%$ in ATE and $16.7\%$ in RTE over RoNIN ResNet. These results quantitatively confirm the significant gains in localization accuracy achieved by our method.
Figure 8: Visualization of sample trajectories across five datasets, comparing MambaIO with Tartan IMU.
To provide a visual assessment of localization performance, we present representative trajectories from each dataset (Fig. 4). For simple strATEht-line or gently curved trajectories (Fig. 4 (a), (c), (d) and (c)), both Tartan IMU and MambaIO closely follow the ground-truth paths. In contrast, for trajectories with multiple turns or complex dynamics (Fig. 4 (b) and (f)), Tartan IMU exhibits noticeable deviations, whereas MambaIO maintains high accuracy. This improvement is attributed to the Laplacian pyramid’s multi-scale decomposition of IMU signals, which enables the model to jointly capture local details and global trends, resulting in a more robust and precise motion representation.
IV-D Ablation Study
To evaluate the effectiveness of the proposed components, we construct two ablation variants: (1) Conv-only, which processes raw IMU signals using only the multi-path convolution module; and (2) Mamba-only, which processes raw IMU signals using only the Mamba module. Table II presents the performance of these variants and the full MambaIO model across the five datasets.
The results indicate that while Conv-only achieves basic localization capability, its accuracy is substantially lower than that of the full MambaIO, suggesting that local modeling alone is insufficient to capture the complex multi-scale dynamics in IMU signals. Similarly, Mamba-only underperforms, highlighting the limited ability of purely global modeling to represent high-frequency local details such as gait transients or turning jitter. Notably, the full MambaIO achieves the best performance across all datasets, significantly surpassing current state-of-the-art methods, which strongly validates the complementary benefits of the high/low-frequency decomposition strategy and the dual-path collaborative architecture.
V Conclusion
This work systematically investigates the impact of coordinate frame selection on pedestrian IO and reveals the advantages of the global frame in motion representation under non-rigid carrier scenarios. Building upon this insight, we propose MambaIO, which decomposes IMU signals into high- and low-frequency components using a Laplacian pyramid and jointly models local details and global motion trends through a multi-path convolution module and a Mamba module. However, the high-frequency component still contains interference noise that is difficult to isolate precisely, and effective separation mechanisms remain lacking—a key direction for future research.